# Repeating Flares, X-ray Outbursts and Delayed Infrared Emission: A Comprehensive Compilation of Optical Tidal Disruption Events

### v1

@@ A Comprehensive Compilation of Optical Tidal Disruption Events
Tidal disruption events (TDEs) have been proposed as valuable laboratories for studying dormant black holes. However, progress in this field has been hampered by the limited number of observed events. In this work, we present TDECat, a comprehensive catalogue of 134 confirmed TDEs (131 optical TDEs and 3 jetted TDEs) discovered up to the end of 2024, accompanied by multi-wavelength photometry (X-ray, ultraviolet (UV), optical, and infrared) and publicly available spectra. We also study the statistical properties, spectral classifications, and multi-band variability of these events. Using a Bayesian Blocks algorithm, we determine the duration, rise time (), decay time (), and their ratio for 103 flares in our sample. We find that these timescales follow a log-normal distribution. Furthermore, our spectral analysis shows that most optical TDEs belong to the TDE-H+He class, followed by the TDE-H, TDE-He, and TDE-featureless classes, which is consistent with expectations from main sequence star disruption. Using archival observations, we identify four new potentially repeating TDEs, namely AT 2024pvu, AT 2022exr, AT2021uvz, and AT 2019teq, increasing the number of known repeating events. In both newly identified and previously known cases, the secondary flares exhibit a similar shape to the primary. We also examine the infrared and X-ray emission from the TDEs in our catalogue, and find that 14 out of the 18 infrared events have associated X-ray emission, strongly suggesting a potential correlation. Finally, we find that for three subsamples (repeating flares, infrared-emitting events, and X-ray-emitting events), the spectral classes are unlikely to be randomly distributed, suggesting a connection between spectral characteristics and multi-wavelength emission. TDEcat enables large-scale population studies across wavelengths and spectral classes, providing essential tools for navigating the data-rich era of upcoming surveys such as the Legacy Survey of Space and Time.
@@ 1 Introduction
A tidal disruption event (TDE) occurs when a star ventures too close to a supermassive black hole (SMBH), where differential gravitational forces exceed the star’s self gravity, causing it to be torn apart by tidal forces. This occurs only if the tidal radius, , which is the critical radius within which an object is disrupted by the SMBH, is larger than the event horizon. There is a theoretical upper limit to the SMBH’s mass known as the Hills mass, since . For a solar mass star, this limit is approximately (Hills 1975). About half of the stellar material remains bound to the SMBH forming an accretion disc, while the other half is ejected (Rees 1988). This process produces a characteristic flare with a sharp rise in brightness, followed by a gradual decay leading to a plateau. These flares can be observed across the electromagnetic spectrum, from radio to hard X-rays and typically exhibit broad hydrogen and helium lines. The expected detection rate of optical TDEs has been estimated at a few galaxy-1 yr-1 (Yao et al. 2023). Over the past half decade, approximately 10-20 new optical TDEs have been detected annually.
@@
The first TDE was identified during the ROSAT all-sky X-ray survey, with the observed soft X-ray emission attributed to a newly formed accretion disc (Bade et al. 1996; Grupe et al. 1999; Saxton et al. 2020). Since then, numerous TDEs have been detected by various instruments and surveys. In the X-ray band, both XMM-Newton and eROSITA have observed multiple TDEs and candidates. The Chandra X-ray Observatory and the BAT telescope aboard the Neil Gehrels Swift Observatory (Roming et al. 2005; Burrows et al. 2005) have played important roles in follow-up X-ray observations of TDEs discovered at optical wavelengths. Additionally, Swift’s UVOT telescope has enabled observations of these events in the ultraviolet.
@@
Optical surveys have significantly contributed to our understanding of TDEs. The All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017; Hart et al. 2023), a global network of small automated telescopes scanning the entire sky for supernovae and other transients, has identified numerous TDEs through their distinctive optical flares. Similarly, the Zwicky Transient Facility (ZTF; Graham et al. 2019; Bellm et al. 2019; Masci et al. 2019), with its high-cadence, wide-field survey capabilities, has played a key role in the discovery of optical TDEs, by rapidly surveying vast areas of the sky. The Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018; Heinze et al. 2018; Smith et al. 2020; Shingles et al. 2021) has also been crucial to long term monitoring of TDEs by providing photometric data in two optical wavebands.
@@
The Gaia photometric Science Alerts system (GSA111http://gsaweb.ast.cam.ac.uk/alerts Hodgkin et al. 2021) has also contributed to TDE discoveries by detecting and providing optical photometry of transient events. Several other surveys were instrumental in early TDE observations, including the Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), Lincoln Near-Earth Asteroid Research (LINEAR; Stokes et al. 2000), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al. 2016; Waters et al. 2020; Magnier et al. 2020a, c, b; Flewelling et al. 2020), the Palomar Transient Factory (PTF Law et al. 2009; Rau et al. 2009) and its successor, the Intermediate Palomar Transient Factory (iPTF222https://www.ptf.caltech.edu/iptf/).
@@
Beyond optical wavelengths, TDEs have been observed in the infrared, initially by the Wide-field Infrared Survey Explorer (WISE; Wright et al. 2010) and later by the Near-Earth Object WISE Reactivation (NEOWISE; Mainzer et al. 2011, 2014). Spectroscopic follow-up observations, including the Sloan Digital Sky Survey (SDSS; York et al. 2000) and the Palomar 200-inch Hale Telescope, have been crucial in classifying TDEs based on their distinctive spectral signatures.
@@
Despite the growing numbers of TDEs no comprehensive catalogue currently exists. Such a resource would enable robust statistical analyses by providing a dataset large enough to conduct meaningful studies, even if the sample is incomplete. The primary goal of this paper is to compile and present all publicly available photometric and spectroscopic data for confirmed optical TDEs up to the end of 2024. Additionally, we investigate different sub-categories of optical TDEs, including those displaying repeating flares, X-ray outbursts, and infrared emission.
@@
Our paper is organized as follows: In Sect. 2, we outline the construction of our main sample of confirmed optical TDEs, along with a sample of TDE candidates. Section 3 describes the data collection process for the catalogue, while Sect. 4 delves into the properties of our sample. Finally, in Sect. 5 we discuss our results and their implications, followed by a summary in Sect. 6.
@@ 2 Sample selection
Our sample consists of events identified either in the Transient Name Server (TNS333https://www.wis-tns.org/; 105 sources) or in the literature (29 sources). We differentiate between confirmed TDEs and TDE candidates by organizing them into two separate catalogues. The main catalogue consists of TDEs that can be found at least in one photometric survey and have an available classification spectrum from the time of the flare. We also include TDEs with featureless spectra (see Sect. 4.2) and the three widely accepted jetted TDEs from the literature. TDEs that do not meet these criteria are designated as TDE candidates and are included in a separate TDE candidates catalogue.
@@ 2.1 Main catalogue
Another TDE sample is presented in Yao et al. (2023) which lists 33 TDEs (see their Table 3). Of these, only 3 are absent from both the Hammerstein et al. (2023) sample and the TNS-TDE sample, namely: AT 2019baf, AT 2019cmw and AT 2020abri. While these 3 were classified as TDEs in Yao et al. (2023), there are uncertainties regarding their classification. AT 2019baf also appears in Somalwar et al. (2023a) as a TDE but it lacks a classification spectrum from the time of the flare. Additionally, Somalwar et al. (2023a) cite an unpublished work as the basis for its classification. A similar situation applies to AT 2019cmw, which is also referenced in an unpublished paper in Yao et al. (2023). AT 2020abri, meanwhile, has an optical spectrum taken 395 days after the peak of its optical flare, well after the event had likely faded. Its classification is based on: 1) persistent blue colour and lack of cooling, which is inconsistent with the majority of SNe and 2) the combination of weak H emission and strong H absorption, which indicate that the host galaxy is post-starburst (where TDE rates are enhanced; see Sect. 4.3). Since none of these three sources have spectra from the time of their flares, we exclude them from the main sample and place them in the TDE candidates sample.
Additionally, several studies focus on individual TDEs that are neither classified in TNS nor included in large sample studies (e.g., van Velzen et al. 2021; Hammerstein et al. 2023; Yao et al. 2023). We include 17 such sources in our catalogue, with detailed descriptions provided in Appendix A.
In total our main catalogue consists of 134 TDEs (131 optical TDEs and 3 jetted TDEs), including all confirmed events up to 2024. As it was briefly mentioned in the Introduction, the creation of a catalogue of all the known optical TDEs so far can allow statistical works, which were previously unable to be carried out due to small sample sizes. We note that this is not a complete sample, since many detected events remain unclassified and certain subtypes (e.g., events detected in infrared) are not fully explored.
@@ 2.2 TDE candidates
Alongside our main TDE sample, numerous transients have been classified as TDE candidates. These sources are included in a supplementary table, available on our GitHub page (see Sect. 3). The setup of this file is shown in Table 1. The first four columns list the name and coordinates, while the fifth and sixth columns include remarks on the candidate status and reference studies, respectively. To compile the TDE candidates sample, we utilized TDExplorer555https://jminding.pythonanywhere.com/main/tdes-by-name, which is a catalogue of TDEs and candidates identified through natural language processing applied to the abstracts of papers.
@@
- | OGLE16aaa | 01:07:20.88 | -64:16:20.70 | TDE candidate in a weak AGN | Wyrzykowski et al. (2017) | ||
@@
- | Swift J1112+82 | 11:11:47.32 | -82:38:44.20 | Likely jetted TDE | Brown et al. (2015) |
@@ 3 Data
The catalogue is available online on a dedicated GitHub666https://github.com/dlangis/TDECat page. On this page we have compiled all publicly available photometric and spectroscopic data for the TDEs in our sample. The full catalogue is also available via a local Python-based app. Below we summarize how and from where we collected the photometric and spectroscopic data included in the catalogue.
@@ 3.1 Optical and Infrared Photometry
For compiling the optical and infrared photometric data, we used the Black Hole Target Observation Manager (BHTOM777https://bhtom.space), a web server designed to provide astronomers easy access to astronomical data and a network of telescopes. One of the key features of BHTOM allows the user to compile all the available, archived photometric data in an interactive plot and a .csv downloadable file. The catalogue includes data from several surveys, namely LINEAR, CRTS, ZTF, iPTF, SDSS, ASAS-SN, ATLAS, NEOWISE, Pan-STARRS and GSA.
We also manually searched for CRTS light curves of LSQ12dyw from the Catalina Surveys Data Release 2888http://nesssi.cacr.caltech.edu/DataRelease/ (Drake et al. 2009). Furthermore, for specific TDEs, we obtained archival photometric data from previously published studies:
Since different surveys use different photometric methods, we converted all magnitudes and estimated flux densities in the AB system (Oke & Gunn 1983) throughout this work. The monochromatic AB magnitude is defined as , where is the spectral flux density and is the zero-point. ZTF magnitudes are calibrated using a source with colour g - r = 0 in the AB photometric system. ATLAS, iPTF, Pan-STARRS and ASAS-SN band photometry is provided in the AB system. On the contrary, CRTS, ASAS-SN band, Swift and NEOWISE999Infrared zero-points for conversion; https://wise2.ipac.caltech.edu/docs/release/allsky/expsup/sec4_4h.html photometry is given in the Vega system. For Vega-to-AB magnitude conversions, we used the values presented in Table 1 of Blanton & Roweis (2007). The Gaia survey uses an internal calibration different from both the AB and Vega systems.
@@ 3.2 X-ray Photometry
To study the X-ray emission from TDE sources we used data from Swift, Chandra and XMM-Newton archives. For each TDE source, we analysed all observations available from these missions, and evaluated X-ray fluxes and spectra for each observation. In addition to Swift-XRT, Chanra-ACIS and XMM-Newton-EPIC data, we searched for X-ray counterparts in the 13th data release of the fourth XMM-Newton serendipitous source catalogue (4XMM-DR13, Webb et al. 2020) and eROSITA main catalogue (eRASS1, Merloni et al. 2024). A detailed description of the X-ray data reduction is provided in Appendix B.

### v2

@@
Tidal disruption events (TDEs) have been proposed as valuable laboratories for studying dormant black holes. However, progress in this field has been hampered by the limited number of observed events. In this work, we present TDECat, a comprehensive catalogue of 134 confirmed TDEs (131 optical TDEs and three jetted TDEs) discovered up to the end of 2024, accompanied by multi-wavelength photometry (X-ray, UV, optical, and infrared) and publicly available spectra. We also study the statistical properties, spectral classifications, and multi-band variability of these events. Using a Bayesian Blocks algorithm, we determined the duration, rise time (), decay time (), and their ratio for 103 flares in our sample. We find that these timescales follow a log-normal distribution. Furthermore, our spectral analysis shows that most optical TDEs belong to the TDE-H+He class, followed by the TDE-H, TDE-He, and TDE-featureless classes, which is consistent with expectations from main-sequence star disruption. Using archival observations, we identified three new potentially repeating TDEs, namely, AT 2024pvu, AT 2022exr, and AT 2021uvz, increasing the number of known repeating events. In both newly identified and previously known cases, the secondary flares exhibit a similar shape to the primary. We also examined the infrared and X-ray emission from the TDEs in our catalogue, and find that 14 out of the 18 infrared events have associated X-ray emission, strongly suggesting a potential correlation. Finally, we find that for three sub-samples (repeating flares, infrared-emitting events, and X-ray-emitting events), the spectral classes are unlikely to be randomly distributed, suggesting a connection between spectral characteristics and multi-wavelength emission. TDEcat enables large-scale population studies across wavelengths and spectral classes, providing essential tools for navigating the data-rich era of upcoming surveys such as the Legacy Survey of Space and Time.
@@ 1 Introduction
A tidal disruption event (TDE) occurs when a star ventures too close to a supermassive black hole (SMBH), where differential gravitational forces exceed the star’s self gravity, causing it to be torn apart by tidal forces. This occurs only if the tidal radius, , which is the critical radius within which an object is disrupted by the SMBH, is larger than the event horizon. There is a theoretical upper limit to the SMBH’s mass known as the Hills mass, since . For a solar mass star, this limit is approximately (Hills 1975). About half of the stellar material remains bound to the SMBH forming an accretion disc, while the other half is ejected (Rees 1988). This process produces a characteristic flare with a sharp rise in brightness, followed by a gradual decay leading to a plateau. These flares can be observed across the electromagnetic spectrum, from radio to hard X-rays and typically exhibit broad hydrogen and helium lines. The expected detection rate of optical TDEs has been estimated at a few galaxy-1 yr-1 (Yao et al. 2023), which is at least around one order of magnitude smaller than the theoretically derived rates for TDEs (e.g. Magorrian and Tremaine 1999; Wang and Merritt 2004; Stone and Metzger 2016). Over the past half decade, approximately 10-20 new optical TDEs have been detected annually.
@@
The first TDE was identified during the ROSAT all-sky X-ray survey, with the observed soft X-ray emission attributed to a newly formed accretion disc (Bade et al. 1996; Grupe et al. 1999; Saxton et al. 2020). Since then, numerous TDEs have been detected by various instruments and surveys. In the X-ray band, both XMM-Newton (e.g. Esquej et al. 2007; Lin et al. 2011; Saxton et al. 2019) and eROSITA (e.g. Saxton et al. 2020; Sazonov et al. 2021; Grotova et al. 2025) have observed multiple TDEs and candidates. The Chandra X-ray Observatory and the BAT telescope aboard the Neil Gehrels Swift Observatory (Roming et al. 2005; Burrows et al. 2005) have played important roles in follow-up X-ray observations of TDEs discovered at optical wavelengths. Additionally, Swift’s UVOT telescope has enabled observations of these events in the ultraviolet.
@@
Optical surveys have significantly contributed to our understanding of TDEs. The All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017; Hart et al. 2023), a global network of small automated telescopes scanning the entire sky for supernovae and other transients, has identified numerous TDEs through their distinctive optical flares (e.g. Holoien et al. 2014, 2016b). Similarly, the Zwicky Transient Facility (ZTF; Graham et al. 2019; Bellm et al. 2019; Masci et al. 2019), with its high-cadence, wide-field survey capabilities, has played a key role in the discovery of optical TDEs (e.g. van Velzen et al. 2021; Hammerstein et al. 2023), by rapidly surveying vast areas of the sky. The Asteroid Terrestrial-impact Last Alert System (ATLAS; Tonry et al. 2018; Heinze et al. 2018; Smith et al. 2020; Shingles et al. 2021) has also been crucial to long-term monitoring of TDEs (e.g. Nicholl et al. 2020; Earl et al. 2025) by providing photometric data in two optical wavebands.
@@
The Gaia photometric Science Alerts system (GSA111http://gsaweb.ast.cam.ac.uk/alerts; Hodgkin et al. 2021) has also contributed to TDE discoveries by detecting and providing optical photometry of transient events (e.g. Wevers et al. 2022). Several other surveys were instrumental in early TDE observations, including the Catalina Real-Time Transient Survey (CRTS; Drake et al. 2009), Lincoln Near-Earth Asteroid Research (LINEAR; Stokes et al. 2000), Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; Chambers et al. 2016; Waters et al. 2020; Magnier et al. 2020a, b, c; Flewelling et al. 2020), the Palomar Transient Factory (PTF; Law et al. 2009; Rau et al. 2009; Arcavi et al. 2014) and its successor, the Intermediate Palomar Transient Factory (iPTF222https://www.ptf.caltech.edu/iptf/; Hung et al. 2017; Blagorodnova et al. 2017, 2019).
@@
Beyond optical wavelengths, TDEs have been observed in the infrared, initially by the Wide-field Infrared Survey Explorer (WISE; Wright et al. 2010) and later by the Near-Earth Object WISE Reactivation (NEOWISE; Mainzer et al. 2011, 2014; Jiang et al. 2016; Dou et al. 2016). Spectroscopic follow-up observations, including the Sloan Digital Sky Survey (SDSS; York et al. 2000; van Velzen et al. 2011) and the Palomar 200-inch Hale Telescope, have been crucial in classifying TDEs based on their distinctive spectral signatures.
@@
Despite the growing numbers of TDEs no comprehensive catalogue currently exists. Such a resource would enable robust statistical analyses by providing a dataset large enough to conduct meaningful studies, even if the sample is incomplete. The primary goal of this paper is to compile and present all publicly available photometric and spectroscopic data for confirmed optical TDEs up to the end of 2024. Additionally, we investigate different sub-categories of optical TDEs, including those displaying repeating flares, X-ray outbursts, and IR emission.
@@
Our paper is organised as follows: In Sect. 2, we outline the construction of our main sample of confirmed optical TDEs, along with a sample of TDE candidates. In Sect. 3 we describe the data collection process for the catalogue, while in Sect. 4 we delve into the properties of our sample. Finally, in Sect. 5 we discuss our results and their implications, and we follow it with a summary in Sect. 6.
@@ 2 Sample selection
Our sample consists of events identified either in the Transient Name Server (TNS333https://www.wis-tns.org/; 105 sources) or in the literature (29 sources). We differentiate between confirmed TDEs and TDE candidates by organising them into two separate catalogues. The main catalogue consists of TDEs that can be found at least in one optical photometric survey and have an available classification spectrum from the time of the flare. We also include TDEs with featureless spectra (see Sect. 4.2) and the three widely accepted on-axis jetted TDEs from the literature. TDEs that do not meet these criteria are designated as TDE candidates and are included in a separate TDE candidates catalogue.
@@ 2.1 Main catalogue
Another TDE sample is presented in Yao et al. (2023) which lists 33 TDEs (see their Table 3). Of these, only three are absent from both the Hammerstein et al. (2023) sample and the TNS-TDE sample, namely, AT 2019baf, AT 2019cmw and AT 2020abri. While these three were classified as TDEs in Yao et al. (2023), there are uncertainties regarding their classification. AT 2019baf also appears in Somalwar et al. (2025a) as a TDE, but it lacks a classification spectrum from the time of the flare. Additionally, Somalwar et al. (2025a) cite an unpublished work as the basis for its classification. A similar situation applies to AT 2019cmw, which is also referenced in an unpublished paper in Yao et al. (2023). AT 2020abri, meanwhile, has an optical spectrum taken 395 days after the peak of its optical flare, well after the event had likely faded. Its classification is based on 1) persistent blue colour and lack of cooling, which is inconsistent with the majority of supernoave and 2) the combination of weak H emission and strong H absorption, which indicate that the host galaxy is post-starburst (where TDE rates are enhanced; see Sect. 4.3). Since none of these three sources have spectra from the time of their flares, we exclude them from the main sample and place them in the TDE candidates sample.
Additionally, several studies focus on individual TDEs that are neither classified in TNS nor included in large sample studies (e.g. van Velzen et al. 2021; Hammerstein et al. 2023; Yao et al. 2023). We include 17 such sources in our catalogue, with detailed descriptions provided in Appendix A.
In total our main catalogue consists of 134 TDEs (131 optical TDEs and three jetted TDEs), including all confirmed events up to 2024. As it was briefly mentioned in the Introduction, the creation of a catalogue of all the known optical TDEs so far can allow for statistical works, which were previously unable to be carried out due to small sample sizes. We note that this is not a complete sample, since many detected events remain unclassified and certain subtypes (e.g. events detected in IR) are not fully explored.
@@ 2.2 TDE candidates
Alongside our main TDE sample, numerous transients have been classified as TDE candidates. These sources are included in a supplementary table, available on our GitHub page (see Sect. 3). We note that this is not an exhaustive list of candidates (e.g. we have not included the IR candidate TDEs from Masterson et al. 2024) and the sources included are not used in any analysis throughout this work. The setup of this file is shown in Table 1. The first four columns list the name and coordinates, while the fifth and sixth columns include remarks on the candidate status and reference studies, respectively. To compile the TDE candidates sample, along with our literature search, we utilised TDExplorer555https://jminding.pythonanywhere.com/main/tdes-by-name, which is a catalogue of TDEs and candidates identified through natural language processing applied to the abstracts of papers.
@@
| - | OGLE16aaa | 01:07:20.88 | -64:16:20.70 | TDE candidate in a weak AGN | Wyrzykowski et al. (2017) | ||
@@
| - | Swift J1112+82 | 11:11:47.32 | -82:38:44.20 | Likely jetted TDE | Brown et al. (2015) |
@@ 3 Data
The catalogue is available online on a dedicated GitHub777https://github.com/dlangis/TDECat page. On this page we have compiled all publicly available photometric and spectroscopic data for the TDEs in our sample. The full catalogue is also available via a local Python-based app. Below we summarise how and from where we collected the photometric and spectroscopic data included in the catalogue. We note that some artifacts could be present in the optical light curves (e.g. see light curve of AT 2019teq) included in the TDECat GitHub page.
@@ 3.1 Optical and infrared photometry
For compiling the optical and IR photometric data, we used the Black Hole Target Observation Manager (BHTOM888https://bhtom.space), a web server designed to provide astronomers easy access to astronomical data and a network of telescopes. One of the key features of BHTOM allows the user to compile all the available, archived photometric data in an interactive plot and a .csv downloadable file. The catalogue includes data from several surveys, namely LINEAR, CRTS, ZTF, iPTF, SDSS, ASAS-SN, ATLAS, NEOWISE, Pan-STARRS and GSA.
We also manually searched for CRTS light curves of LSQ12dyw from the Catalina Surveys Data Release 2999http://nesssi.cacr.caltech.edu/DataRelease/ (Drake et al. 2009). Furthermore, for specific TDEs, we obtained archival photometric data from previously published studies:
Since different surveys use different photometric methods, we converted all magnitudes and estimated flux densities in the AB system (Oke and Gunn 1983) throughout this work. The monochromatic AB magnitude is defined as , where is the spectral flux density and is the zero point. ZTF magnitudes are calibrated using a source with colour g - r = 0 in the AB photometric system. ATLAS, iPTF, Pan-STARRS and ASAS-SN band photometry is provided in the AB system. On the contrary, CRTS, ASAS-SN band, Swift and NEOWISE101010IR zero points for conversion; https://wise2.ipac.caltech.edu/docs/release/allsky/expsup/sec4_4h.html photometry is given in the Vega system. For Vega-to-AB magnitude conversions, we used the values presented in Table 1 of Blanton and Roweis (2007). The Gaia survey uses an internal calibration different from both the AB and Vega systems.
@@ 3.2 X-ray photometry
To study the X-ray emission from TDE sources we used data from Swift, Chandra, and XMM-Newton archives. For each TDE source, we analysed all observations available from these missions, and evaluated X-ray fluxes and spectra for each observation. In addition to Swift-XRT, Chanra-ACIS and XMM-Newton-EPIC data, we searched for X-ray counterparts in the 13th data release of the fourth XMM-Newton serendipitous source catalogue (4XMM-DR13, Webb et al. 2020) and eROSITA main catalogue (eRASS1, Merloni et al. 2024). A detailed description of the X-ray data reduction is provided in Appendix B.